Shift summary for DAY shift 9/Jan/2000


Subject: Shift summary for DAY shift 9/Jan/2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:11:46 CST


                     Shift summary for day shift 9/Jan/2000

On Shift: PG, AA

Hours of beam available at all intensities: 6.5
Hours of beam available at nominal intensity: 6.5
Hours of E799 data (special runs not included): 6.5
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

A TeV magnet quench. Starting and Stopping runs, KFREND's.
========================================================================

Tapes brought to FCC this shift:

None
========================================================================

Accesses during shift: Two

Jason and Philippe twice to swap an ET controller.
Ashkan and Philippe to check on the bad DCOR2Y power supply.
========================================================================

Runs ended this shift:

15445 - 6.3M - BEAM_799
15446 - 1.7M - BEAM_799
 
=======================================================================

Overview of shift:

E799 data-taking.

========================================================================

Description of Shift:

run spill time comment
------------------------------------------------------------------------

15445 8 07:15 ktevana error followed by pipeline errors. Reset
                       the pipeline and KFREND.

          21 07:41 Trigger hangs, KFREND.

        69-71 08:37 Etotal temperature error, hcc error and trigger
                       hangs. Reload ETOT maps + a number of KFREND's.
    
                08:45 The ET temp error persists, page Jason H.

       121-123 09:48 Trigger hangs, KFREND.

                10:00 Onsplit tape UPM056 mounted on ktev1.

           141 10:13 Quench at D3, est 1hr. Page people on the list.
          
           147 10:22 Stop the run. Update Ped and Csi constants.

15446 1 11:06 Start the run, no beam.

                11:24 JH and PG go on access to replace ET controller #6
                       in BW crate #4. But JH unhappy with the result,
                       they go to swap it back.
                       AA and PG on access to check on bad DCOR PS
                       channel. Turned out to be a monitoring problem and
                       got fixed.

                11:45 LB swaps an ADC module.

            48 12:11 Beam returns.

         65-68 12:35 Trigger hangs, KFREND's.

            70 12:39 ktevana error, KFREND.
          
            87 13:02 Stop the run due to L3 accept fraction errors.

15447 1 13:12 Start the run.

                14:11 onsplit tape UPF076 mounted on ktev4.

            71 14:48 TRD plane 2 tripped off.

========================================================================
Items to Follow up from previous Shift

  DCOR2Y LV power supply for DC2 read out is fixed.

  Philippe labeled 28 QKE tapes for short reg run and other special runs

========================================================================
Items to Follow up on Next Shift

  Don't forget to hit "reset quartz" button after each KFREND.

  If beam goes away for more than an hour (non-owl) page people listed on
  page 214 of the logbook.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:11:49 CST